OpenStudy (anonymous):

A 1.20-g sample of propane (C3H8) is burned in a bomb calorimeter. The temperature of the calorimeter rises from 20.0°C to 30.0°C. The calorimeter has a mass of 2.000 kg and a specific heat of 2.95 J/(g • °C). If the molar mass of propane is 44.0 g/mol, what is the molar heat of combustion of propane? 6,490 kJ/mol 4,330 kJ/mol 2,160 kJ/mol 59.0 kJ

OpenStudy (aaronq):

First find the thermal energy absorbed by the water: \(\sf q_{absorbed}=m_{water}*C_p*\Delta T\) then scale it to "per mole" with the mass given to get the molar heat of combustion \(\sf H^o_{comb}\). \(\sf q_{absorbed}=-q_{released}=- H^o_{comb}*n_{propane}\) (n=moles)

OpenStudy (anonymous):

I have no clue where to start. I don't even see water in the question.

OpenStudy (anonymous):

I have 11 Minutes to find out how to do this to get the answer.

OpenStudy (jfraser):

water is used in the calorimeter to absorb the energy and increase in temperature. That's how we measure the energy released by the reaction

OpenStudy (anonymous):

so there is 2kg of water.

OpenStudy (aaronq):

They have integrated the mass of the water with the calorimeter. The first equation should be: \(q_{abs}=m_{cal}*C^{cal}_p*\Delta T\)

OpenStudy (anonymous):

But i don't understand how i would take the numbers and plug them into the equation.

OpenStudy (aaronq):

m=mass, \(C_P\)=specific heat capacity, \(\Delta T=T_f-T_i\) (temp change)

OpenStudy (anonymous):

10 C is ΔT

OpenStudy (aaronq):

yep

OpenStudy (anonymous):

2.95 is Cp

OpenStudy (anonymous):

mass would be 1.20 g?

OpenStudy (anonymous):

Would it be the mass of the sample or water.

OpenStudy (aaronq):

nope, thats the mass of the propane (which is releasing the heat), you would use the mass ABSORBING the heat, which is the mass of the calorimeter, 2kg.

OpenStudy (aaronq):

look at the units though, heat capacity is given in grams, so you need to convert

OpenStudy (anonymous):

111Mol?

OpenStudy (aaronq):

you're using mass here, not moles

OpenStudy (anonymous):

Oh! then it would be 2000 grams

OpenStudy (aaronq):

yep!, now multiply those values together to get q

OpenStudy (anonymous):

59000

OpenStudy (aaronq):

dont forget the units. Now we wanna use the other equation to scale the heat released to a molar value. We need the moles of propane, use \(\sf moles=\dfrac{mass}{Molar~mass}\)

OpenStudy (anonymous):

so 59000/44

OpenStudy (anonymous):

= 1337

OpenStudy (aaronq):

not quite. You didnt find the moles of propane, you just divided the energy by the molar mass. \(\sf H^o_{comb}=\dfrac{q}{\dfrac{mass}{Molarmass}}=\dfrac{q*Molar~ mass}{mass}\)

OpenStudy (anonymous):

.027

OpenStudy (anonymous):

Mol of propane

OpenStudy (aaronq):

okay, now divide the heat released by the moles

OpenStudy (anonymous):

so 59000 divided by .027

OpenStudy (aaronq):

yep

OpenStudy (aaronq):

you should always write units

OpenStudy (anonymous):

Yeah, sorry, i got a hugggggggeeeeee number.

OpenStudy (aaronq):

it should be pretty large

OpenStudy (anonymous):

2,185,185.185

OpenStudy (aaronq):

right, now convert it to kJ, 1000 J = 1 kJ

OpenStudy (anonymous):

Then divide by 1000

OpenStudy (aaronq):

yes

OpenStudy (anonymous):

so it would be 2185.2 which is closest to C.

OpenStudy (anonymous):

Kj/m

OpenStudy (aaronq):

yep. remember to write mole, as "mol" not just "m"

OpenStudy (anonymous):

Oops yeah, thank you!

OpenStudy (aaronq):

no problem!